OpenStudy (anonymous):

use the echelon method to solve system of equalities X+y=10 5x-y=44

OpenStudy (shayaan_mustafa):

hello tishagwen. are you there?

OpenStudy (shayaan_mustafa):

hello mr.

OpenStudy (anonymous):

sorry yea im here

OpenStudy (anonymous):

and its mrs.. ")

OpenStudy (shayaan_mustafa):

oh .. hehehehhehe... mrs. or right. let us solve it. Well I don't know how to use arrays on this site. so kindly. try to understand me.

OpenStudy (anonymous):

the problem changed sorry

OpenStudy (shayaan_mustafa):

yes. what is it?

OpenStudy (anonymous):

X+y=10 5x-y=44

OpenStudy (shayaan_mustafa):

miss it is the same question that you've posted.

OpenStudy (anonymous):

if i understand right.. ill need to take the first set and multiply by 5 and the the 2nd one multiply by-1 right?

OpenStudy (shayaan_mustafa):

no no. here you are solving the equation with Gauss elimination method(Echelon form).

OpenStudy (anonymous):

oh ok

OpenStudy (shayaan_mustafa):

so first write equation as matrices form. 1 1 10 5 -1 44 OK?

OpenStudy (anonymous):

got it

OpenStudy (shayaan_mustafa):

ok now Replace row 2 (i.e. R2) with the row operation R2 new=-4*R1+R2 in order to convert some elements in the row into the desired value 0.

OpenStudy (shayaan_mustafa):

do this step. then tell me.

OpenStudy (anonymous):

x-5y=-34

OpenStudy (shayaan_mustafa):

no no my dear.

OpenStudy (shayaan_mustafa):

do the row operation on R2(row 2)

OpenStudy (anonymous):

what do you mean

OpenStudy (shayaan_mustafa):

means. multiply R1 with 4 and minus it from R2 i.e. R2new=R2-4R1 Does this make sense to you?

OpenStudy (anonymous):

not really i know nothing of the matrix stuff

OpenStudy (shayaan_mustafa):

now look at pic and here I am going to solve it for you. keep the pic in mind. OK

OpenStudy (anonymous):

how did u get 4 i thought it was 5

OpenStudy (shayaan_mustafa):

yes yes it is 5 sorry. it is a mistake.

OpenStudy (anonymous):

ok

OpenStudy (shayaan_mustafa):

step#2) multiply R1 with 5